OpenStudy (anonymous):

if f(x)=log2 (x+4), what is f^-1(3)?

OpenStudy (anonymous):

be clarify very well

ganeshie8 (ganeshie8):

solve below :- 3 = log2 ( x+4 )

OpenStudy (mathmale):

ganeshie8's reply can be solved for x by raising 2 to the power 3 on the left side of the equation and to the power log2 (x+4) on the right side: 2^3 = 2^log2(s+4) Since y=log2 x and y=2^x are inverses of each other, 2^log2(x+4) simplifies to x + 4. Thus, 8 = x + 4. x + ?

OpenStudy (mathmale):

Another, more roundabout, approach would be to actually find the inverse of y = base2log (x+4): a) replace f(x) with y b) interchange x and y (replace x with y and y with x) c) solve the resulting equation for y d) replace this y with -1 f (x) -1 e) The result is f (x) = 2^x -4.. f) evaluate this at x=3.